A CD starts from rest and accelerates to an angular frequency of 3.5 rev/s. Calculate the disc s average period T and centripetal velocity at the edge of the disc when radius is 2.0*10^-2m.


1
Expert's answer
2022-05-02T18:05:54-0400

"T=\frac 1\nu=0.29~s,"

"v=\omega r=2\pi \nu r=0.44~\frac ms."